OpenStudy (anonymous):

solve for x cos(5x)+cos(2x)=0

OpenStudy (anonymous):

use cosA+cosB identity

OpenStudy (anonymous):

\[x = 2 (\pi n-\tan^(-1)(\sqrt(5/3-(2 (1+i \sqrt(3)))/(3 (1/14 (13+3 i \sqrt(3)))^(1/3))-1/3 2^(2/3) (1-i \sqrt(3)) (1/7 (13+3 i \sqrt(3)))^(1/3))))\]

OpenStudy (anonymous):

i only know the identity of cos(A+B)

OpenStudy (anonymous):

tell me what cos(A+B)+cos(A-B) is

OpenStudy (anonymous):

so you mean B=0?

OpenStudy (anonymous):

nope just tell me what you get by using the formulae

OpenStudy (anonymous):

cos(A)cos(B)-sin(A)sin(B)+cos(A)cos(B)+sin(A)sin(B)

OpenStudy (anonymous):

=2cos(A)cos(B) right?

OpenStudy (anonymous):

=> cos(5x)cos(2x)=0

OpenStudy (anonymous):

no dont jump to conclusions now put A+B=x and A-B=y

OpenStudy (anonymous):

which means A=(x+y)/2 and B=(x-y)/2

OpenStudy (anonymous):

now, what you get by cosx+cos y

OpenStudy (anonymous):

why need the set A and x?

OpenStudy (anonymous):

i am trying to get you to the formula. this x is not the x in the problem. instead its a general variable.

OpenStudy (anonymous):

would you mind to give me all steps and lead it to the answer Pi/7 + 4pi/7 k, 3pi/7+ 4pi/7 k, pi/3 + 4pi/3 k

OpenStudy (anonymous):

x=pi/3